OpenStudy (anonymous):

What is the value of 1 over 2^-2 x^-3 y^5 for x = 2 and y = -4?

OpenStudy (anonymous):

substitute

OpenStudy (anonymous):

I don't know how to do this at all..

OpenStudy (anonymous):

ok look

OpenStudy (anonymous):

you have this 2^-2 x^-3 y^5 now x= 2 and y = -4 2^-2 (2)^-3 (-4)^5 =1/4. 1/8. (-1024) = -1024/32 do it